2 moles of an ideal gas are at a temperature of -13degrees Celsius and a pressure of 2atm. What volume in dm3 will the gas occupy at that temperature?
[ R= 0.082]
1
Expert's answer
2020-04-23T13:05:45-0400

PV = nRT

V = nRT / P = (2 × 0.082 × (273.15 - 13)) / 2 = 21.33 dm3